A gene called “PRNP” produces the prion protein. In humans there are two alleles for the gene,
known as “Met” and “Val”. People who are homozygous for the Met allele can contract
“variant CJD”, the fatal human form of mad cow disease associated with the consumption of
infected beef.
In Denmark, a survey of 352 individuals found that 131 were homozygous for Met, 168 were
heterozygous, and 53 were homozygous for Val. Is the PRNP gene in Hardy-Weinberg
equilibrium? Show your calculations.
1
Expert's answer
2016-10-14T13:04:03-0400
The frequency of allele for Met is (2*131+168)/2*352=0.61

The frequency of allele for Val is (2*53+168)/2*352=0.39

According to the calculated frequencies, the number of individuals, who are homozygous for Met, should be: 0.61^2 *352=131

In the same way, the expected number of heterozygous individuals is 2*0.61*0.39*352=168

Similarly, the expected number of individuals who are homozygous for Val is: 0.39^2*352=53

Thus, the PRNP gene is in Hardy-Weinberg equilibrium